A continuous r.v. X has the p.d.f. as follows: f(x) = Zx(9 – x2 ),​0 ≤ x ≤ 3
= 0,​otherwise

Calculate the value of Z?
1
Expert's answer
2020-09-13T18:20:53-0400

solution


The PDF of X is


"f(x) = Zx(9-x^2)" Where "0 \leq x \leq3"


The sum of all probabilities i.e the integral of the PDF for all values of X "=1"



"1=\textstyle\int_0^3 Zx(9-x^2)"

"\frac{1}{Z}=\textstyle\int_0^3 (9x-x^3)"

"\frac{1}{Z}= (4.5x^2-0.25x^4) \textstyle\mid_0^3"

"\frac{1}{Z}= 4.5(3)^2-0.25(3)^4-0"

"\frac{1}{Z} = 20.25"

"Z= \frac{4}{81} = 0.0494"

Answer: Z is 0.0494